# Scaling Large Motion Models with Million-Level Human Motions

### v2

@@ 5.3 Effectiveness of Motionbook
Motion Feature Design. We conduct an ablation study to evaluate the impact of different motion feature design, including H3D-Format, SMPL-D130, SMPL-D135, SMPL-D263, and SMPL-D268 (detailed in Appendix A.3). As shown in Table LABEL:tab:motion_feature, retrieval performance remains consistent across all features, with no significant differences. Similarly, quantitative evaluations of rendered samples reveal no noticeable variation among different features. This indicates that none of the tested features provide a clear performance advantage. However, we observe that H3D-Format discards original motion information, requiring additional time-consuming processing to restore missing details — an issue that complicates downstream applications. Considering simplicity, computation cost, and information completeness, we finally select SMPL-D135 as the motion feature.
@@ 6 Conclusion
In this paper, we explore how to advance the field of large motion model. To this end, we introduce MotionLib, the first million-level dataset comprising over 1.2 million motions with hierarchical texts. Building on MotionLib, we present key insights into scaling up both data and model size for large-scale motion training. Furthermore, we propose MotionBook, a novel motion encoding approach designed to maximize the benefits when trained on extensive motion data. MotionBook incorporates compact yet lossless features to represent motion, and introduces a novel 2D-LFQ motion quantization method that treats each motion sequence as a 2D image, constructing a finite-scale codebook that eliminates the need for token lookups. Leveraging these advancements, we train Being-M0, a large motion model that achieves SoTA results compared to current counterparts.
@@ A.2 Implementation Details
For a fair comparison with previous works, we implement oue model Being-M0 based on two varionts of motion tokenizers: one with a vector quantized (VQ) codebook and another with a 2D lookup-free (2D-LFQ) codebook. By default, our Being-M0 is trained using 2D-LFQ. For the motion tokenizer, we implement the VQ codebook $\mathbbm{C}\in\mathbbm{R}^{1024\times 512}$ with an embedding dimensionality of $d=512$. The resulting discrete codes are incorporated as additional vocabulary for the LLM. As a comparison, the LFQ codebook has a size of $163842^{16}=$ 16384. The motion encoder $\mathbbm{E}$ uses a temporal downsampling rate of $\alpha=4$. We experiment with four large language model (LLM) architectures to construct our large motion model: GPT2-medium (Radford et al., 2019), LLaMA2-7b, LLaMA2-13b (Touvron et al., 2023), and LLaMA3.1-8b (Dubey et al., 2024). The motion tokenizer is trained with a learning rate of 1e-4 and a batch size of 256 for 300K iterations. For training the large motion model, full parameter tuning is performed on 8$\times$×A800 GPUs with a batch size of 1024 over 300 epochs. The learning rate is set to 2e-4 for GPT2-medium and 2e-5 for the LLaMA models. Implementation on more LLMs is on the way.
@@ B.2 Motion Data Refinement
To ensure physically plausible motion data, we refine the extracted motion using an RL-based policy (Luo et al., 2023). Before this refinement step, we enhance motion quality following the methodology of Lin et al. (2024). We begin by estimating 2D human keypoints and their confidence scores using the pretrained VitPose model (Xu et al., 2022), then infer 3D keypoints via another pretrained 3D keypoint estimation model (Sárándi et al., 2023). To improve stability and consistency, we apply temporal smoothing and enforce 3D bone length constraints during triangulation. As previously mentioned, we fit the SMPL-X body model (Pavlakos et al., 2019) to each frame in MotionLib using the approach from Shin et al. (2024), followed by iterative optimization to align model parameters with the estimated 2D and 3D joint positions. We further refine global motion and camera poses using a global motion optimization technique based on Yuan et al. (2022), ensuring consistency with the original video evidence. For motion sequences affected by noise or occlusion, we employ RoHM (Zhang et al., 2024a) to reconstruct complete and physically plausible motions.
@@ C.1 Ablation of Large Motion Model Training
C.1.1 Hierarchical vs. Single-level Text.
To investigate the effectiveness of hierarchical descriptions, we conduct ablation experiments, as shown in Table LABEL:tab:app_hiera_basic. We compare the training results using hierarchical descriptions, which include body-level and part-level labels, against those using only whole-body description. The results indicate that hierarchical text significantly enhances the model’s semantic understanding, leading to improved semantic matching of generated motions.
TRAIN TEXT R@1 $\uparrow$ R@3 $\uparrow$ FID $\downarrow$ MMDist $\downarrow$
Real 0.297 0.634 0.004 2.068
single-level 0.162 0.371 7.018 4.682
hierarchical 0.166 0.375 6.936 4.484
C.1.2 LoRA vs. Full Parameter Fine-tuning
@@
C.1.3 Learning from Scratch vs. Fine-tuning
@@
C.1.4 With vs. Without Description Mask
@@
C.1.5 Encoder-decoder vs. Decoder-only
@@
C.1.6 Slow convergence of large motion models

### v3

@@ 5.3 Effectiveness of Motionbook
Motion Feature Design. We conduct an ablation study to evaluate the impact of different motion feature design, including H3D-Format, SMPL-D130, SMPL-D135, SMPL-D263, and SMPL-D268 (detailed in Appendix A.3). As shown in Table LABEL:tab:motion_feature, retrieval performance remains consistent across all features, with no significant differences. Similarly, quantitative evaluations of rendered samples reveal no noticeable variation among different features. This indicates that none of the tested features provide a clear performance advantage. However, we observe that H3D-Format discards some original motion information, as its rotations are obtained through inverse kinematics. This requires additional time-consuming processing to restore missing details — an issue that complicates downstream applications. Considering simplicity, computation cost, and information completeness, we finally select SMPL-D135 as the motion feature.
@@ 6 Conclusion
In this paper, we explore how to advance the field of large motion model. To address the issue of data scarcity in this domain, we introduce MotionLib, the first million-level dataset comprising over 1.2 million motions with hierarchical texts. Building on MotionLib, we present key insights into scaling up both data and model size for large-scale motion training. Furthermore, we propose MotionBook, a novel motion encoding approach designed to maximize the benefits when trained on extensive motion data. MotionBook incorporates compact yet lossless features to represent motion, and introduces a novel 2D-LFQ motion quantization method that treats each motion sequence as a 2D image, constructing a finite-scale codebook that eliminates the need for token lookups. Leveraging these advancements, we train Being-M0, a large motion model that achieves SoTA results compared to current counterparts.
Acknowledgements
This work was supported by NSFC under Grant 62450001. The authors would like to thank the anonymous reviewers for their valuable comments and advice.
@@ A.2 Implementation Details
For a fair comparison with previous works, we implement oue model Being-M0 based on two varionts of motion tokenizers: one with a vector quantized (VQ) codebook and another with a 2D lookup-free (2D-LFQ) codebook. By default, our Being-M0 is trained using 2D-LFQ. For the motion tokenizer, we implement the VQ codebook $\mathbbm{C}\in\mathbbm{R}^{1024\times 512}$ with an embedding dimensionality of $d=512$. The resulting discrete codes are incorporated as additional vocabulary for the LLM. As a comparison, the LFQ codebook has a size of $163842^{16}=$ 16384. The motion encoder $\mathbbm{E}$ uses a temporal downsampling rate of $\alpha=4$. We experiment with four large language model (LLM) architectures to construct our large motion model: GPT2-medium (Radford et al., 2019), LLaMA2-7b, LLaMA2-13b (Touvron et al., 2023), and LLaMA3.1-8b (Dubey et al., 2024). The motion tokenizer is trained with a learning rate of 1e-4 and a batch size of 256 for 300K iterations. For training the large motion model, full parameter tuning is performed on 8$\times$×A800 GPUs with a batch size of 1024 over 100 epochs. The learning rate is set to 2e-4 for GPT2-medium and 2e-5 for the LLaMA models.
@@ B.2 Motion Data Refinement
To ensure physically plausible motion data, we refine the extracted motion using an RL-based policy (Luo et al., 2023). Before this refinement step, we enhance motion quality following the methodology of Lin et al. (2024). We begin by estimating 2D human keypoints and their confidence scores using the pretrained VitPose model (Xu et al., 2022), then infer 3D keypoints via another pretrained 3D keypoint estimation model (Sárándi et al., 2023). To improve stability and consistency, we apply temporal smoothing and enforce 3D bone length constraints during triangulation. As previously mentioned, we fit the SMPL-X body model (Pavlakos et al., 2019) to each frame in MotionLib using the approach from Shin et al. (2024), followed by iterative optimization to align model parameters with the estimated 2D and 3D joint positions. We further refine global motion and camera poses using a global motion optimization technique based on Yuan et al. (2022), ensuring consistency with the original video evidence. For motion sequences affected by noise or occlusion, we employ RoHM (Zhang et al., 2024b) to reconstruct complete and physically plausible motions.
@@ C.1 Ablation of Large Motion Model Training
C.1.1 LoRA vs. Full Parameter Fine-tuning
@@
C.1.2 Learning from Scratch vs. Fine-tuning
@@
C.1.3 With vs. Without Description Mask
@@
C.1.4 Encoder-decoder vs. Decoder-only
@@
C.1.5 Slow convergence of large motion models